The product of two numbers is 24 and their quotient is 2/3. The smaller of the two numbers is?
Answer by lwsshak3(11628)

The product of two numbers is 24 and their quotient is 2/3. The smaller of the two numbers is?
***
let x=larger of two numbers
let y=smaller of two numbers
xy=24
x=24/y
y/x=2/3
2x=3y
48/y
3y^2=48
y^2=16
y=4
x=24/y=6
The smaller of the two numbers is 4